# Reconstruction the scalar-torsion gravity version from the frame of exact cosmological solutions

### v3

@@ Abstract
We consider the cosmological models based on the scalar-torsion gravity implying non-minimal coupling between torsion and the scalar field as modification of teleparallel equivalent of general relativity. Based on generalized exact solutions of the equations of cosmological dynamics, the type of scalar-torsion gravity was reconstructed. Also, based on observational constraints on the values of the parameters of cosmological perturbations, the type of the coupling function was determined. It was noted that any models of cosmological inflation obtained on the basis of the proposed approach can be verified by observational constraints.
1 Introduction
@@
Currently, a description of the inflationary stage of the evolution of the universe with including some inflaton fields for the case of general relativity and its various modifications (with and without inflaton fields) is necessary for constructing correct cosmological models [1, 2, 3]. The first works on inflationary cosmology [4, 5, 6, 7] taught their further development, and at the moment there are a large number of cosmological models describing both stage of accelerated expansion of the universe with different types of the scalar fields based on the Einstein gravity theory [1, 2], the teleparallel equivalent of the general relativity (TEGR) [8, 9] and their various extensions and modifications [3, 10, 11, 12, 13, 14, 15, 16, 17] are actively used.
Also, various mechanisms of the particle production are considered, including rapid oscillations of the inflaton near the minimum potential in accordance with the reheating process [18], instant preheating [19], curvaton reheating [20] and gravitational particle production [21], including the case of non-minimum coupling of torsion and material fields [22] as well.
Thus, at the moment, there is no specific theory of gravity or cosmological model that would be considered to unconditionally describe the evolution of the universe, and different cosmological models corresponding to observational data can be considered as actual ones.
The construction and analysis of cosmological models can be considered in two aspects, namely by the analysis of the possible physical effects induced by the different versions of the gravity theories [1, 2, 3], and, also, by the reconstruction of the model’s parameters from the various observational constraints including the results of measurements of anisotropy and polarization of CMB [23, 24, 25].
An important method for analyzing cosmological models is the use of the slow-roll approximation approach, which takes into account that the early universe at the inflationary stage expands close to exponential expansion law or close to the de Sitter stage [1, 2, 3]. However, exact cosmological solutions provide a clearer understanding of the specifics of inflationary processes or the parameters of cosmological models [26]. Methods for constructing exact cosmological solutions based on the different symmetries of cosmological dynamic equations, the nonlocal conservation laws, various form-invariant transformations and many other methods can be used to construct and analyze exact cosmological solutions (see, for example, in [2]). The background dynamics of the expansion of the universe and the evolution of cosmological perturbations can be considered on the basis of both exact and approximate cosmological solutions [1, 2, 3].
According to the inflationary paradigm, quantum fluctuations of the scalar field generate corresponding perturbations of the space-time metric, which are the source of large-scale structure and relict gravitational waves [1, 2, 3].
As a basis for analyzing the compliance of cosmological models with observational constraints on the values of the parameters of cosmological perturbations, the dependence of the tenor-to-scalar ratio on the spectral index of scalar perturbations $r=r(1-n_{S})$ is usually used to verify the cosmological models by observational constraints following from the results of measurements of anisotropy and polarization of CMB [27, 28].
@@
At the moment, a large number of cosmological models corresponding to the observational constraints on the parameters of cosmological perturbations are considered [3, 11, 12, 13, 14, 15, 16]. For cosmological models based on general relativity (or TEGR), agreement with observational data will lead to an unambiguous determination of the model parameters including the potential of a scalar field [1, 2]. However, for the case of cosmological models based on modified theories of gravity, agreement with observational data can be obtained for ambiguously defined parameters [29, 30, 31, 32, 33, 34].
@@
For example, in the case of cosmological models based on the scalar-tensor gravity theories [34, 35, 36, 37, 38, 39] in paper [34] it was shown that these models can meet observational constraints not due to the choice of model parameters, but due to a specific relationship between the parameters for arbitrary cosmological models.
Thus, within the framework of the approach proposed in [34], any cosmological model can be verified using observational data after finding the appropriate connection between the parameters of the modified theory of gravity and the parameters of the material source of gravity, which is considered to be scalar fields. As a criterion for finding such connections, the exact solutions of the cosmological dynamics equations and known types of scalar-tensor gravity and scalar field potentials were used [34].
@@
The description of the teleparallel gravity, its modifications including the non-minimal coupling of torsion and the scalar field, and their applications in cosmology were previously considered in many works (see, for example, in [40, 41, 42, 43]).
@@
The properties of the cosmological models with scalar-torsion gravity expressed in a fairly general form $f(T,\phi)_{STG}=f(T,\phi)$ were considered earlier in [44, 45, 46] based on the slow-roll approximation. Also, specific cosmological models based on the exact solutions of dynamic equations were considered in [47].
This paper presents new generalized exact solutions of the cosmological dynamic equations for the certain type of scalar-torsion gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$. Also, the relationship between the potential of the scalar field and the type of the coupling function between the scalar field and torsion was reconstructed. The exact solutions obtained are generalized both for arbitrary scalar field potentials (positive and negative due to the the signature reversal symmetry) [48, 49, 50, 51, 52, 53, 54], for arbitrary cosmological dynamics and for arbitrary evolution of scalar fields as well.
The first property of the cosmological models based on the proposed version of the scalar-torsion gravity is the possibility of their verification for an arbitrary model’s parameters for the case of a linear dependence of the tensor-to-scalar ratio on the spectral index of scalar perturbations $r\sim(1-n_{S})$, in contrast to the case of Einstein gravity (or TEGR) for which models with the dependence $r\sim(1-n_{S})$ are unverified by observational constraints [1, 2].
The second property of the proposed version of the scalar-torsion gravity is that relation $r\sim(1-n_{S})$ is preserved for an arbitrary type of cosmological dynamics, which implies an arbitrary choice of a scale factor $a=a(t)$ or Hubble parameter $H=H(t)$ for cosmological models verified by these observational constraints.
[…]
Thus, on the one hand, the proposed approach makes it possible to construct a large number of phenomenologically correct cosmological models, however, on the other hand, this property was obtained for the specific version of gravity theory, the discussion of which is devoted to this work.
The paper is organized as follows. In Sec. 2, the exact solutions of the cosmological dynamic equations for the scalar-torsion gravity of type $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$ are obtained, where $T$ is the torsion scalar, and function $G(\phi)$ defines the character of the non-minimal coupling between a scalar field and torsion. It is also shown that for the case of the quasi-de Sitter expansion of the universe, this type of the scalar-torsion gravity is close to the case of the teleparallel equivalent of general relativity. In Sec. 3, expressions for the parameters of cosmological perturbations are obtained in explicit form for the inflation based on the scalar-torsion gravity under consideration. Sec. 4 discusses a new classification of cosmological models according to the order of expansion of the dependence of the tensor-to-scalar ratio on the spectral index of scalar perturbations $r=r(1-n_{S})$. Relationships between the Hubble parameter $H$, the non-minimal coupling function $G$ and the scalar field potential $V$ are obtained for cosmological models based on the proposed version of scalar-torsion gravity, which are verified by observational constraints on the values of the parameters cosmological perturbations in the first order of this expansion $r\sim(1-n_{S})$. It is also shown that, taking into account these relationships, cosmological models with arbitrary dynamics of the expansion of the universe (subject to the slow-roll conditions at the inflationary stage) based on the scalar-torsion gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$ will correspond to observational constraints on the values of the parameters of cosmological perturbations. In Sec. 5, as an example of the proposed approach to the construction the cosmological models based on scalar-torsion gravity $f(T,\phi)_{STG}\sim G(\phi)\sqrt{T}$, the case of the exponential scalar field potential is considered. It is shown that for this case it is possible to reconstruct exact expressions for all other parameters of cosmological models. Also, using the examples of well-known cases of exponential power-law inflation [58, 59, 60, 61, 62] and Quasi-Steady State cosmological model [63, 64, 65, 66], it is shown that the proposed gravity theory allows both to describe the first and second accelerated expansion of the universe and to propose a solution to the Hubble tension problem with taking into account the observational constraints on the parameters of cosmological perturbations. Sec. 6 discusses the general properties of cosmological models based on the proposed version of scalar-torsion gravity in comparison with models based on the other modified gravity theories. Our results are summarized in Conclusion.
2 Cosmological models in generalized scalar-torsion gravity
@@
The generalized action for cosmological models based on the scalar-torsion gravity with non-minimal coupling of a scalar field $\phi$ and torsion $T$ in the system of units $c=8\pi G=M^{-2}_{p}=1$ can be written as [44, 45, 46]
@@
which is a proper tetrad naturally associated with the vanishing spin connections $\omega^{A}_{~{}B\mu}=0$, and which leads to the spatially flat Friedmann-Robertson-Walker (FRW) metric [44, 45, 46]
@@
The background equations for this model can be written as [44, 45, 46]
@@
It should be noted, that the torsion scalar is $T=6H^{2}$ for the case of the spatially flat Friedmann-Robertson-Walker metric [44, 45, 46].
@@
The exact solutions of the system of equations (7)–(9) are represented by the special form of the function $f(T,\phi)$, connected with the potential $V(\phi)$ and kinetic term $\frac{1}{2}\omega(\phi)\dot{\phi}^{2}$ and can be written as
@@
$\dot{\phi}=\frac{\sqrt{6}\alpha_{2}}{\alpha_{1}}\frac{V(\phi)}{G_{,\phi}},$ (12)
where $G(\phi)$ is arbitrary differentiable function of a scalar field.
The function
$f(T,\phi)_{STG}=\alpha_{1}G(\phi)\sqrt{T},$ (13)
fixed the version of the scalar-torsion gravity and $\alpha_{1}$, $\alpha_{2}$ are the normalization constants, which allow one to consider the function (10) with different signs.
Using expression $\ddot{\phi}=\frac{d\dot{\phi}}{d\phi}\dot{\phi}$, one can verify by direct substitution, that solutions (10)–(12) satisfy all three dynamic equations (7)–(9).
Let us note, that by using $G_{,\phi}=\dot{G}/\dot{\phi}$ expression (12) can be rewritten as follow
$\dot{G}=\frac{\sqrt{6}\alpha_{2}}{\alpha_{1}}\,V(\phi).$ (14)
@@
As it was noted in [45], for the function
$f(T,\phi)=-\frac{1}{2}T-V(\phi),$ (15)
one has the case of TEGR, and dynamic equations (7)–(9) for the case $\omega=1$ are reduced to well known ones
$3H^{2}=\frac{1}{2}\dot{\phi}^{2}+V(\phi),$ (16)
$\dot{\phi}^{2}=-2\dot{H},$ (17)
$\ddot{\phi}+3H\dot{\phi}+V_{,\phi}=0,$ (18)
which are the same ones for inflationary models based on the Einstein gravity [1, 2].
@@
Nevertheless, to describe the gravitational interaction we will consider the function (13) in general form as exact solutions consequence of cosmological dynamics equations. Besides, the form of a coupling function $G=G(\phi)$ in expression (13) is not determined.
Now, based on the nature of early universe accelerated expansion (close to exponential one) and observational constraints on the values of cosmological perturbations parameters, we reconstruct the connection between the coupling function $G=G(\phi)$ and the scalar field potential $V=V(\phi)$. For this purpose, let us consider cosmological perturbations in inflationary models based on the scalar-torsion gravity and generalized exact solutions (10)–(14) of cosmological dynamic equations (7)–(9).
3 Cosmological perturbations
In accordance with the theory of cosmological perturbations, quantum fluctuations of the scalar field induce the corresponding perturbations of the space-time metric during the inflationary stage. In the linear order of cosmological perturbation theory, the observed anisotropy and polarization of cosmic microwave background radiation (CMB) [1, 2] are explained by the influence of two types of perturbations, namely, scalar and tensor ones.
Observational constraints on the parameters of cosmological perturbations due to the modern observations of the anisotropy and polarisation of CMB are [27, 28]
@@ 3 Cosmological perturbations
$n_{S}=0.9663\pm 0.0041,$ (22)
$r<0.032.$ (23)
Cosmological perturbations for the inflationary models based on the scalar-torsion gravity and action (1) were considered on the basis of the Arnowitt-Deser-Misner (ADM) decomposition of the tetrad field in order to define the perturbed metrics in [45]. Further, based on a standard analysis of the equations for the evolution of cosmological perturbations, expressions for the parameters of cosmological perturbations were obtained (see, for details, in [45]).
The expressions for the parameters of cosmological perturbations at the crossing of the Hubble radius ($k=aH$) for the cosmological models based on action (1) can be written as follows [45]
$\mathcal{P}_{S}=\frac{H^{2}}{8\pi^{2}Q_{S}}\left[1+2\eta_{\mathcal{R}}\ln\left(\frac{k}{aH}\right)\right]_{k=aH}=\frac{H^{2}}{8\pi^{2}Q_{S}},$ (24)
$\mathcal{P}_{T}=\frac{H^{2}}{2\pi^{2}Q_{T}},$ (25)
$n_{S}-1=-2\epsilon-\eta+2\eta_{\mathcal{R}},$ (26)
$n_{T}=-2\epsilon-\delta_{f_{,T}},$ (27)
$r=\frac{\mathcal{P}_{T}}{\mathcal{P}_{S}}=16\delta_{\omega X},$ (28)
where $\mathcal{P}_{S}$ and $\mathcal{P}_{T}$ are the power spectra of scalar and tensor perturbations, $n_{S}$ and $n_{T}$ are the spectral index of the scalar and tensor perturbations and $r$ is the tensor-to-scalar ratio, also, $n_{T}<0$ corresponds to the red tilt tensor spectrum, $n_{T}>0$ corresponds to the blue tilt tensor spectrum and $n_{T}=0$ corresponds to the flat tensor spectrum.
Also, the parameters $Q_{S}$ and $Q_{T}$ are [45]
@@
where all these parameters much less than unity ($\epsilon\ll 1$) due to the quasi de Sitter nature of the inflationary stage of the evolution of the early universe.
On the basis of expressions (11)–(12) and (14) for the proposed exact solutions we obtain following expression
@@
Thus, form expression (34) we have the following conditions: $\alpha_{2}>0$ for the positive potentials $V>0$ and $\alpha_{2}<0$ for the negative potentials $V<0$ of a scalar field.
The examples of different cosmological models with positive and negative potentials of a scalar field one can find in [48, 49, 50, 51, 52, 53, 54].
@@
Thus, we can write the spectral indices of the scalar and tensor perturbations and the tensor-to-scalar ratio as
@@
4 Reconstruction of non-minimal coupling function from linear dependence $r\sim(1-n_{S})$
@@
Since, the zeroth order term in this expansion $r(0)=\beta_{0}=0$ from condition $r(n_{S}=1)=0$ corresponding to the flat Harrison-Zel’dovich spectrum [67, 68], we can rewrite expression (41) in the following form
@@
Since the first term in expression (42) makes the main contribution to the value of the tensor-scalar ratio, we can consider the dependence $r=r(1-n_{S})$ with a reasonable degree of accuracy in the first order on the basis of expressions (38)–(39).
Thus, we can consider the new specific classification of cosmological models (regardless of gravity type or cosmological model’s parameters) according to non-zero orders of expansion of dependence $r=r(1-n_{S})$ as follows: first-order models for $\beta_{1}\neq 0$, second-order models for $\beta_{1}=0$ and $\beta_{2}\neq 0$ etc.
For the purpose to obtain the connection between parameters of the cosmological models under consideration for the first-order models in explicit form with arbitrary Hubble parameter (i.e. for arbitrary cosmological dynamics), we consider the special case $\alpha_{1}=-1$, and from expressions (38)–(39) we obtain
@@
$r=-16\frac{\dot{G}}{HG}.$ (44)
The linear dependence between tensor-to-scalar ratio and spectral index of scalar perturbations we define as follows
$r\simeq\frac{8}{m}(1-n_{S}),$ (45)
@@
Also, we note that dependence (45) is explicit one and it doesn’t depend on the background dynamic parameters, i.e. this dependence is valid for any parameters of the cosmological model, including the times corresponding to the e-folds number $N\approx 60$ at the end of the inflationary stage [1, 2].
Also, on the basis of the observational constraints $n_{S}\simeq 0.97$ and $r<0.032$ from (45) for verifiable inflationary models one has following constraint $m>7.5$.
Figure 1: The dependence $r=r(n_{S})$ for different values of the parameter $m=7,8,12,24$ with constraints on the tensor-to-scalar ratio $r$ due to the Planck TT,TE,EE+lowE+lensing+BAO+BICEP2/Keck Array observations [27, 28].
Figure 1 shows dependence (45) for various values of $m$. It can be seen that under condition $m>7.5$, the values of the spectral index of scalar perturbations $n_{S}$ and the tensor-scalar ratio $r$ fall into the internal region of allowed values, which corresponds to the possibility of verifying the inflationary model according to the observational constraints.
@@
From equations (14) and (46) for such a models we finally obtain
@@
for any type of the scalar field evolution $\phi=\phi(t)$, where $G_{0}$ is a some arbitrary constant.
@@
Therefore, taking into account quasi de Sitter regime of accelerated expansion of the early universe $H\simeq const$, we can represent the exact expression for the coupling function as follows
$G(\phi)=const\times\left[V(\phi)\right]^{-\frac{1+k}{2(m-1)}},$ (49)
@@
where parameter $k$ defines the deviations from the pure de Sitter regime.
We also note that expression (49) is one of the possible ways to parameterize deviations from the de Sitter stage in expression for the coupling function, and we use it as a fairly convenient one to analyze exact solutions of the cosmological dynamics equations for the certain inflationary models.

### v4

@@ Abstract
We consider cosmological models based on the scalar-torsion gravity implying non-minimal coupling between torsion and the scalar field with certain relations between model’s parameters. Based on observational constraints on the values of the parameters of cosmological perturbations, the type of the coupling was determined. It was noted that any inflationary models constructed on the basis of the proposed approach can be verified by observational constraints.
I Introduction
@@
Currently, a description of the inflationary stage of the evolution of the universe with including some inflaton fields for the case of general relativity and its various modifications is necessary for constructing correct cosmological models Baumann:2014nda ; Chervon:2019sey ; Odintsov:2023weg . At the moment there are a large number of cosmological models describing both stage of accelerated expansion of the universe with different types of the scalar fields based on the Einstein gravity theory Baumann:2014nda ; Chervon:2019sey , the teleparallel equivalent of the general relativity (TEGR) Aldrovandi:2013wha ; Bahamonde:2021gfp and their various extensions and modifications Cai:2015emx ; Fomin:2020hfh ; Fomin:2022ozv ; Gonzalez-Espinoza:2019ajd ; Gonzalez-Espinoza:2020azh ; Chervon:2023gio are actively used.
@@
For cosmological models based on general relativity (or TEGR), agreement with observational data will lead to an unambiguous determination of the model parameters including the potential of a scalar field Baumann:2014nda ; Chervon:2019sey ; Bahamonde:2021gfp . However, for the case of cosmological models based on modified theories of gravity, agreement with observational data can be obtained for ambiguously defined parameters Cai:2015emx ; Fomin:2020hfh ; Fomin:2022ozv ; Gonzalez-Espinoza:2019ajd ; Gonzalez-Espinoza:2020azh ; Chervon:2023gio .
@@
For example, for the case of cosmological models based on the scalar-tensor gravity theories in paper Fomin:2022ozv it was shown that these models can satisfy observational constraints not due to the choice of model parameters, but due to a specific relationship between the parameters for arbitrary cosmological models. To obtain such relation, the exact solutions of the cosmological dynamics equations for known types of scalar-tensor gravity were used Fomin:2022ozv .
@@
The properties of the cosmological models with scalar-torsion gravity expressed in a fairly general form were considered earlier in Gonzalez-Espinoza:2019ajd ; Gonzalez-Espinoza:2020azh based on the slow-roll approximation. Also, specific cosmological models based on the exact solutions of dynamic equations were considered in Chervon:2023gio .
In this case, we consider generalized exact solutions of the cosmological dynamic equations for inflation based on the scalar-torsion gravity for arbitrary Hubble parameter to reconstruct a class of verifiable inflationary models. Also, the relationship between the potential of the scalar field and the type of the scalar-torsion gravity was reconstructed in explicit form.
II Cosmological models in generalized scalar-torsion gravity
@@
The generalized action for cosmological models based on the scalar-torsion gravity with non-minimal coupling of a scalar field $\phi$ and torsion $T$ in the system of units $c=8\pi G=M^{-2}_{p}=1$ can be written as Gonzalez-Espinoza:2019ajd ; Gonzalez-Espinoza:2020azh
@@
which is a proper tetrad naturally associated with the vanishing spin connections $\omega^{A}_{~{}B\mu}=0$, and which leads to the spatially flat Friedmann-Robertson-Walker (FRW) metric Gonzalez-Espinoza:2019ajd ; Gonzalez-Espinoza:2020azh
@@
The background equations for this model can be written as Gonzalez-Espinoza:2019ajd ; Gonzalez-Espinoza:2020azh
@@
It should be noted, that the torsion scalar is $T=6H^{2}$ for the case of the spatially flat Friedmann-Robertson-Walker metric Gonzalez-Espinoza:2019ajd ; Gonzalez-Espinoza:2020azh .
@@
The exact solutions of the system of equations (7)–(9) for the special form of the function $f(T,\phi)$ can be written as follow
@@
$\dot{G}=\dot{\phi}\,G_{,\phi}=\frac{\sqrt{6}\alpha_{2}}{\alpha_{1}}V(\phi),$ (12)
where function $f(T,\phi)_{STG}=\alpha_{1}G(\phi)\sqrt{T}$ fixed the version of the scalar-torsion gravity, $G=G(\phi)$ is arbitrary differentiable function of a scalar field, and $\alpha_{1}$, $\alpha_{2}$ are the normalization constants.
Using expression $\ddot{\phi}=\frac{d\dot{\phi}}{d\phi}\dot{\phi}$, one can verify by direct substitution, that relations (10)–(12) satisfy all three dynamic equations (7)–(9).
@@
Also, we note, that for the case teleparallel equivalent of general relativity (TERG) corresponding function is Bahamonde:2021gfp
$f(T,\phi)=-\frac{1}{2}T-V(\phi).$ (13)
@@
Now, we reconstruct the connection between the coupling function $G=G(\phi)$ and the scalar field potential $V=V(\phi)$ based on the nature of early universe accelerated expansion (close to exponential one) and observational constraints on the values of cosmological perturbations parameters. For this purpose, let us consider cosmological perturbations in inflationary models based on the scalar-torsion gravity and generalized exact solutions (10)–(12) of cosmological dynamic equations (7)–(9).
III Cosmological perturbations
In accordance with the theory of cosmological perturbations, quantum fluctuations of the scalar field induce the corresponding perturbations of the space-time metric during the inflationary stage. In the linear order of cosmological perturbation theory, the observed anisotropy and polarization of cosmic microwave background radiation (CMB) Baumann:2014nda ; Chervon:2019sey are explained by the influence of two types of perturbations, namely, scalar and tensor ones.
Observational constraints on the parameters of cosmological perturbations due to the modern observations of the anisotropy and polarisation of CMB are Planck:2018vyg ; Galloni:2022mok
@@ III Cosmological perturbations
$n_{S}=0.9649\pm 0.0042,$ (17)
$r<0.028.$ (18)
The expressions for the parameters of cosmological perturbations at the crossing of the Hubble radius ($k=aH$) for the cosmological models based on action (1) can be written as follows Gonzalez-Espinoza:2020azh
$\mathcal{P}_{S}=\frac{H^{2}}{8\pi^{2}Q_{S}}\left[1+2\eta_{\mathcal{R}}\ln\left(\frac{k}{aH}\right)\right]_{k=aH}=\frac{H^{2}}{8\pi^{2}Q_{S}},~{}~{}~{}~{}\mathcal{P}_{T}=\frac{H^{2}}{2\pi^{2}Q_{T}},$ (19)
$n_{S}-1=-2\epsilon-\eta+2\eta_{\mathcal{R}},~{}~{}~{}~{}n_{T}=-2\epsilon-\delta_{f_{,T}},~{}~{}~{}~{}r=\frac{\mathcal{P}_{T}}{\mathcal{P}_{S}}=16\delta_{\omega X},$ (20)
where $\mathcal{P}_{S}$ and $\mathcal{P}_{T}$ are the power spectra of scalar and tensor perturbations, $n_{S}$ and $n_{T}$ are the spectral index of the scalar and tensor perturbations and $r$ is the tensor-to-scalar ratio. Also, $n_{T}<0$ corresponds to the red tilt tensor spectrum, $n_{T}>0$ corresponds to the blue tilt tensor spectrum and $n_{T}=0$ corresponds to the flat tensor spectrum.
Also, the parameters $Q_{S}$ and $Q_{T}$ are Gonzalez-Espinoza:2020azh
@@
where all these parameters much less than unity due to the quasi de Sitter inflationary regime of the early universe’s dynamics Gonzalez-Espinoza:2019ajd ; Gonzalez-Espinoza:2020azh .
On the basis of expressions (11)–(12) for the proposed exact solutions we obtain following expression
@@
Thus, we can write the spectral indices of the scalar and tensor perturbations and the tensor-to-scalar ratio at the crossing of the Hubble radius as
@@
IV Reconstruction of the parameters of verified inflationary models
@@
Since, the zeroth order term in this expansion $r(0)=\beta_{0}=0$ from condition $r(n_{S}=1)=0$ corresponding to the flat Harrison-Zel’dovich spectrum Baumann:2014nda ; Chervon:2019sey , we can rewrite expression (33) in the following form
@@
Thus, we can consider the new specific classification of cosmological models (regardless of gravity type or cosmological model’s parameters) according to the orders of expansion of dependence $r=r(1-n_{S})$.
Since the first term in expression (34) makes the main contribution to the value of the tensor-to-scalar ratio, we can consider the dependence $r=r(1-n_{S})$ with a reasonable degree of accuracy in the first order $r\sim(1-n_{S})$ on the basis of expressions (30)–(31).
For the purpose to obtain the connection between parameters of the cosmological models under consideration for the first-order models with $r\sim(1-n_{S})$ in explicit form with arbitrary Hubble parameter (i.e. for arbitrary cosmological dynamics), we consider $\alpha_{1}=-1$. For this case, from expressions (30)–(32) we obtain
@@
$r=-16\frac{\dot{G}}{HG},$ (36)
$n_{T}=-3\epsilon+\frac{r}{16}.$ (37)
Also, we note, that for the partial case $G\sim H$, from (35)–(37) we obtain
$n_{S}-1=-4\epsilon+2\delta,~{}~{}~{}n_{T}=-2\epsilon,~{}~{}~{}~{}r=16\epsilon,$ (38)
where $\delta=-\frac{\ddot{H}}{2H\dot{H}}$, that corresponds to the case of TEGR Gonzalez-Espinoza:2019ajd for the scalar-torsion gravity $f(T,\phi)_{STG}=-G(\phi)\sqrt{T}$ under consideration. In such a case, inflationary models with only some certain potentials of the scalar field will satisfy the observational constraints on the parameters of cosmological perturbations Chervon:2023gio .
In general case, the linear dependence between tensor-to-scalar ratio and spectral index of scalar perturbations can be defined as follows
$r=\frac{8}{m}(1-n_{S}),$ (39)
@@
Also, on the basis of the observational constraints (17)–(18) from (39) for verifiable inflationary models one has following constraint $m>11$. We also note that any future constraints on the values of the cosmological perturbation parameters only lead to a change in the estimate of the constant parameter $m$. Thus, models of cosmological inflation based on relation (39) are fundamentally verified by observational constraints following from observations of the anisotropy and polarization of CMB Planck:2018vyg ; Galloni:2022mok .
@@
From equations (12) and (40) for such a models we obtain
@@
for any type of the scalar field evolution $\phi=\phi(t)$.
@@
Therefore, taking into account quasi de Sitter regime of accelerated expansion of the early universe $H\simeq const$, we can represent the expression for the coupling function as follows
$G(\phi)\sim\left[V(\phi)\right]^{-\frac{1+k}{2(m-1)}},$ (43)
@@
where parameter $|k|\ll 1$ defines the deviations from the pure de Sitter stage.